OpenStudy (anonymous):

Differentiate f(x) = 30 + 10e^-0.05x

OpenStudy (anonymous):

\[f(x) = 30 + 10e ^{-0.05x}\]

OpenStudy (anonymous):

well, the derivative of e^ax is just a*e^ax the derivative of a constant is 0 and the 10 in front of e^x can be pulled out of the derivative, and put back in afterwards. thus, it'll be -.5e^-.05x

OpenStudy (anonymous):

\[f'(x) = 10e ^{^{-0.05x}} + 0\] Do I multiply 10 by 0.05? \[f'(x) = -.5e ^{-0.05x} \]

OpenStudy (anonymous):

yep yep that is the right process you can take the derivative of each term (constant and e^x), the derivative of e^x is pretty much like the power rule, a*e^ax, if a = 1 then e^x = e^x.

OpenStudy (anonymous):

thank you!! :)